# KAG-Thinker: Teaching Large Language Models to Think with Human-like Reasoning Process

### v1

@@ Abstract
In this paper, we introduce KAG-Thinker, a novel human-like reasoning framework built upon a parameter-light large language model (LLM). Our approach enhances the logical coherence and contextual consistency of the thinking process in question-answering (Q&A) tasks on domain-specific knowledge bases (KBs) within LLMs. This framework simulates human cognitive mechanisms for handling complex problems by establishing a structured thinking process. Continuing the Logical Form guided retrieval and reasoning technology route of KAG v0.7, firstly, it decomposes complex questions into independently solvable sub-problems(also referred to as logical forms) through breadth decomposition, each represented in two equivalent forms-natural language and logical function-and further classified as either Knowledge Retrieval or Reasoning Analysis tasks, with dependencies and variables passing explicitly modeled via logical function interfaces. In the solving process, the Retrieval function is used to perform knowledge retrieval tasks, while the Math and Deduce functions are used to perform reasoning analysis tasks. Secondly, it is worth noting that, in the Knowledge Retrieval sub-problem tasks, LLMs and external knowledge sources are regarded as equivalent KBs. We use the knowledge boundary model to determine the optimal source using self-regulatory mechanisms such as confidence calibration and reflective reasoning, and use the depth solving model to enhance the comprehensiveness of knowledge acquisition. Finally, instead of utilizing reinforcement learning, we employ supervised fine-tuning with multi-turn dialogues to align the model with our structured inference paradigm, thereby avoiding excessive reflection. This is supported by a data evaluation framework and iterative corpus synthesis, which facilitate the generation of detailed reasoning trajectories. Experimental results show that KAG-Thinker outperforms state-of-the-art models on seven benchmark datasets, achieving an average improvement of 4.1%. We further demonstrate its practical effectiveness through a medical Q&A system integrated with domain-specific knowledge. We train KAG-Med-Thinker model with 14b-parameter based on synthetic medical corpora using the same approach, respectively, and verify their applicability in specialized applications.
@@ 1 Introduction
Recent advances in large language models (LLMs) OpenAI et al. (2024); Qwen et al. (2025) have shown impressive performance in Q&A tasks. However, their ability to solve multi-hop problems Wei et al. (2022), especially those that require external retrieval of up-to-date information Jin et al. (2025a), has certain limitations. Previous work primarily focused on how to use external retrievers, broadly falling into two categories: (1) decoupling retrieval from training, and (2) autonomously deciding whether to use the retriever during training. The first category primarily involves retrieval augmentation generation (RAG), with representative works including Lewis et al. (2020), Trivedi et al. (2023), and Li et al. (2025). These models often use the LLM’s input to retrieve relevant passages, which are then added to the LLM’s context. This helps the LLM leverage external knowledge when answering questions. The second category mainly concerns reinforcement learning-based retrieval augmentation generation, exemplified by Jin et al. (2025b), Sun et al. (2025), and Wang et al. (2025b). These methods generate a specific token <search> that indicates an external search engine query is required. Once the retriever provides results, these are then fed back into the model as input, allowing it to resume content generation. Although these methods improved LLM utilization of retrievers, they do not truly provide a step-by-step approach for solving complex multi-hop problems. Their approaches are more concerned with how LLMs arrive at the correct answer, without ensuring the rigor, stability, or logical consistency of the underlying problem-solving process. Consequently, these methods are often unusable in high-stakes or sensitive domains such as finance, medicine, and legal applications.
@@
To address these problems and inspired by how humans solve complex problems, we propose an innovative structured human-like cognitive reasoning framework, as shown in Figure 1. We use logical form to decompose complex problems broadly and solve them in depth. Logical form improves the logic, rigor, and stability of the LLM reasoning process. When humans solve complex problems, they only perform external information retrieval on knowledge that they do not know. Similarly to this process, we do not need to perform external information retrieval for all decomposed sub-problems. Therefore, we design an LLM’s knowledge boundary determination module to reduce unnecessary external retrieval and mitigate noise from the retrieved content. If the LLM cannot confidently answer a certain sub-problem using its own knowledge, we rely on external search results. However, search results are often unreliable and noisy. To address this, we created a Focusing-and-Reasoning module that extracts the core relevant information. To facilitate a clearer understanding, we will detail the following key innovations.
@@
Logical Form based Reasoning. Prior RAG approaches Lewis et al. (2020); Trivedi et al. (2023); Li et al. (2025), for both questions and sub-problems, are restricted to plain text representations, thereby preventing them from effectively utilizing high-quality knowledge graphs. To tackle this issue, logical form proposes four operation functions tailored for frequently encountered problems (see Appendix 8). Each operation comprises two components: a natural language segment (Step) and a logical function expression segment (Action). For generic retrievers like E5 and BGE-M3, we can directly employ the content within the logical form’s Step. Conversely, for knowledge graph retrievers, the Action portion of the Logical form can be utilized. Crucially, logical form facilitates the propagation of dependencies among interdependent sub-problems. For instance, consider "Step1: Who is the director of Hit Parade Of 1947? Action1: Retrieval(s=s1:film[‘Hit Parade Of 1947’], p=p1:director, o=o1:director)" followed by "Step2: When did #1 die? Action2: Retrieval(s=o1, p=p2:deathtime, o=o2:deathtime)". The variables and enable the seamless transfer of text and expressions from the outcome of the first planning step. In this manner, the logical form enhances the logic, rigor, and stability of LLM thinking.
@@
Breadth Decomposition and Depth Solving. Direct retrieval often proves insufficient for complex multi-hop questions, failing to provide immediate answers. For example, "Which film has the director who died first, Hit Parade Of 1947 or Khiladi 420?". To address such issues, we have devised a framework based on breadth decomposition and depth solving. At a macroscopic level, the decomposition of complex questions into logical forms is generally straightforward and typically does not necessitate recourse to external knowledge. For instance, the aforementioned example can be broken down into five constituent sub-questions: "Who is the director of Hit Parade Of 1947?", "When did #1 die?", "Who is the director of Khiladi 420?", "When did #3 die?", and "Which film was directed by the director who died first according to #2 and #4?". The resolution of each individual sub-problem, conversely, may entail ranging from to (maximum number of searches) retrieval operations, contingent upon its inherent complexity. This in-depth approach to sub-problem solving enables each sub-problem to assimilate adequate external knowledge, thereby reaching a solvable state.
@@
Knowledge Boundary Determination. Most prior approaches Asai et al. (2023); Gutierrez et al. (2024) initiate retrieval for every question or sub-problem, even when the information resides within the LLM’s inherent knowledge. However, LLMs do not necessitate external lookups for such straightforward queries. To mitigate this redundancy, we design a knowledge boundary determination module that eliminates unnecessary external retrieval operations. Prior to knowledge boundary determination, the LLM first generates answers to the sub-problems. To ensure the precision of this knowledge boundary determination, we employ two strategies: (1) prompt-based confidence assessment and (2) likelihood-based confidence assessment. Only when both conditions are satisfied is considered that the solution to the sub-problem does not require external retrieval, and the answer generated by the LLM is directly adopted.
@@
Ultimately, our research yields a human-like cognitive reasoning framework offering substantial advancements, notably: (1) Facilitating more effective utilization of high-quality KBs, (2) Contributing to a more logically coherent, rigorous, and stable approach to complex problem solving, (3) Enabling autonomous decisions regarding the reliance on intrinsic model knowledge versus external information, thereby reducing unnecessary queries, and (4) Exhibiting a greater capacity to manage noise within retrieved data.
@@ 2 Related Work
Retrieval-Augmented Generation (RAG) integrates relevant, real-time information from external KBs through a retrieval mechanism, thereby enhancing its responses with contextual richness and factual grounding Borgeaud et al. (2022); Dong et al. (2023); Luo et al. (2024). Despite these advantages, a persistent challenge in RAG systems is that retrieved content can often be mixed with irrelevant or erroneous information due to the inherent limitations of current retrieval solutions. To mitigate these challenges, researchers have introduced various components designed to enhance the robustness and effectiveness of RAG systems Gao et al. (2023); Li et al. (2024); Tan et al. (2024). A key focus in recent studies has been on enabling RAG systems to autonomously determine when to leverage retrieval mechanisms. For example, Jin et al. (2025b); Sun et al. (2025); Wang et al. (2025b) introduced a special token to explicitly trigger an external search, with the retrieval results subsequently integrated into the model to support content generation. However, Ren et al. (2025) identified limitations in LLM regarding their ability to accurately distinguish between internal and external knowledge, as well as their difficulty in reconciling conflicts between the two. To address this, they proposed Prior judgment and Poster judgments to evaluate confidence levels and assess response correctness. In parallel, another significant line of research has focused on developing adaptive strategies to optimize how retrieval is utilized, based on query complexity. Jeong et al. (2024) developed a classifier to intelligently decide between iterative, single, or no retrieval strategies depending on the complexity of the query. Building upon this idea, Islam et al. (2024) proposed a hybrid adaptive retrieval method: The model initially generates an answer without retrieval and then dynamically decides the need for retrieval based on its confidence, with the aim of balancing performance gains with inference speed. Furthermore, Tang et al. (2025) extended this work by applying a reinforcement learning-based framework to select the most appropriate retrieval strategy for a given query. While these diverse approaches have significantly advanced the integration and adaptive use of retrieval mechanisms within RAG systems, they currently lack a comprehensive, step-by-step framework for effectively addressing complex multi-hop reasoning problems. This critical area remains an open challenge for further research.
@@ 3 Approach
We propose an innovative problem solving framework that simulates a human-like reasoning process, as shown in Figure 2. At its core, this framework uses logical form to enable both broad problem decomposition and in-depth problem solving. By integrating a knowledge boundary determination module, we minimize unnecessary external queries, balancing the LLM’s internal knowledge with external information. For sub-problems that require external information, a dedicated focusing and reasoning module is employed to reduce noise in the retrieved content and extract essential information.
@@ 3.1.2 Depth Solving
Constrained by sub-problem representations and retriever capabilities, many sub-problems defy resolution through single-turn retrievals. We have engineered a depth-oriented resolution framework for Retrieval sub-problem, as detailed in the right part of Figure 2. During depth solving, we prompt the LLM to iteratively search across multi-level and multi-dimensional contexts. The iterative process continues until one of the following conditions is met: (1) the maximum number of turns is reached, or (2) the model generates a conclusive response enclosed between the designated answer tokens <answer> and </answer>. Before conducting depth solving of the Retrieval sub-problem, we first perform a knowledge boundary determination on the sub-problem to assess whether the LLM can directly answer it. If the LLM can confidently answer the current sub-problem, there is no need for depth solving, and the answer can be directly obtained; otherwise, an depth solving process will be initiated. After retrieving relevant content, we conduct focusing and reasoning analysis, allowing the model to determine contextually whether to initiate the next action. If a subsequent action is required, a new Logical Form is generated; otherwise, the sub-question answer is produced. This process iterates interactively until either the maximum number of turns is reached or the sub-question answer is obtained.
@@ 3.4 Logical Form Solver
Second, as shown in Figure 4, we build a hybrid graph retriever in KAG to retrieve this structured index. This hybrid retrieval method consists mainly of five steps: Step 1. DPR: Use the natural language Step part in the specific logical form to retrieve the top-k chunks by matching the name or content with vector similarity. Step 2. Seed Node Focusing: Identify initial relevant knowledge units, atomic queries, and entities of the sub-problem as seed nodes for Personalized PageRank (PPR). The key steps are shown in Figure 5. Step 3. PPR: The top-k entities, top-k knowledge units, and top-k atomic queries retrieved in Step 2 are integrated as seed nodes, and PPR Yang et al. (2024b) runs on the entire graph. Identify the top-k chunks through probability propagation on the path. Step 4. Rerank: For the chunks obtained by DPR and PPR, top-k chunks are selected, respectively, and then rerank them through Reciprocal Rank Fusion (RRF) Cormack et al. (2009). Step 5. Collect Facts: The top-k chunks and the top-k SPOs matched by the logical form are sent to KAG-Thinker as supporting facts for answer generation in this step.
@@ 4.1 Experimental Settings
Evaluation Metric. While EM Yu et al. (2024) provides strict verification of answer precision; however,its binary nature is overly restrictive for open-ended scenarios where semantically equivalent answers may exist legal phrasing differences. The F1 metric therefore serves as a nuanced supplement, particularly valuable when evaluating free-form textual output. Comparison Methods. To systematically assess KAGThinker’s effectiveness, we establish a three-tiered evaluation framework comprising the following state-of-the-art baselines: (1) Non-Retrieval Paradigms Naive Generation: Direct answer synthesis without external knowledge integration; Chain-of-Thought (CoT): Explicit cognitive pathway formalization through sequential reasoning traces Wei et al. (2022). (2) Retrieval-Augmented Architectures Naive RAG: Standard retrieval-generation pipeline without iterative optimization Lewis et al. (2020); IRCoT: Multi-cycle retrieval-reasoning coordination with dynamic feedback mechanisms Trivedi et al. (2023); Search-o1: Agent-mediated search workflow integration in reasoning processes Li et al. (2025). (3) Reinforcement Learning Frameworks R1-Gen: Pure RL-optimized generation without search engine interfacing DeepSeek-AI et al. (2025); Search-R1: Multimodal trajectory optimization combining search interactions with reasoning steps Jin et al. (2025b); ZeroSearch: Supervised LLM transformation into dual-function retrieval module (relevant/noisy document generation) Sun et al. (2025); StepSearch: PPO-based search LLM training with incremental exploration Wang et al. (2025b).

### v2

# KAG-Thinker: Interactive Thinking and Deep Reasoning in LLMs via Knowledge-Augmented Generation

@@ Abstract
In this paper, we introduce KAG-Thinker, which upgrade KAG to a multi-turn interactive thinking and deep reasoning framework powered by a dedicated parameter-light large language model (LLM). Our approach constructs a structured thinking process for solving complex problems, enhancing the the logical coherence and contextual consistency of the reasoning process in question-answering (Q&A) tasks on domain-specific knowledge bases (KBs) within LLMs. Following the Logical Form guided retrieval and reasoning technology route of KAG, this framework first decomposes complex questions into independently solvable sub-problems (which are also referred to as logical forms) through breadth decomposition. Each such logical form is represented in two equivalent forms—natural language and logical function—and subsequently classified as either a Knowledge Retrieval or Reasoning Analysis task. Dependencies and parameter passing between these tasks are explicitly modeled via logical function interfaces. In the solving process, the Retrieval function performs retrieval tasks. It retrieves one-hop structured and unstructured information of specified knowledge unit. While the Math and Deduce functions are used to perform reasoning analysis tasks. Secondly, it is worth noting that, in the Knowledge Retrieval sub-problem tasks, LLMs and external knowledge sources are regarded as equivalent KBs. We use the knowledge boundary module to determine the optimal source using self-regulatory mechanisms such as confidence calibration and reflective reasoning, and use the depth solving module to enhance the comprehensiveness of knowledge acquisition. Finally, instead of utilizing reinforcement learning, we employ supervised fine-tuning with multi-turn interactive thinking and deep reasoning to align the model with our structured reasoning paradigm, thereby avoiding excessive reflection. This is supported by a data evaluation framework and iterative corpus synthesis, which facilitate the generation of detailed reasoning trajectories. Experimental results show that our model outperforms state-of-the-art models on seven benchmark datasets, achieving an average improvement of 4.1%. In addition, KAG-Thinker not only enhances the reasoning stability of the KAG framework but also achieves performance comparable to the model with 10 times its parameters. We further demonstrate its practical effectiveness through a medical Q&A system integrated with domain-specific knowledge. We train KAG-Med-Thinker model with 14b-parameter based on synthetic medical corpora using the same approach, respectively, and verify their applicability in specialized applications.
@@ 1 Introduction
Recent advances in large language models (LLMs) OpenAI et al. (2024); Qwen et al. (2025) have shown impressive performance in Q&A tasks. However, their ability to solve multi-hop problems Wei et al. (2022), especially those that require external retrieval of up-to-date information Jin et al. (2025a), has certain limitations. Previous work primarily focused on how to use external retrievers, broadly falling into two categories: (1) decoupling retrieval from training, and (2) autonomously deciding whether to use the retriever during training. The first category primarily involves retrieval augmentation generation (RAG) with predefined workflow, the representative works include Lewis et al. (2020), Trivedi et al. (2023), and Li et al. (2025a). These models often use the question to retrieve relevant passages, which are then added to the LLM’s context. This helps the LLM leverage external knowledge when answering questions. The second category mainly concerns to mimic the deeper and more deliberative reasoning as "slow thinking" Pan et al. (2025), a paradigm inspired by human cognitive processes, to determine the timing of retrieval dynamically. Exemplified by Jin et al. (2025b), Sun et al. (2025), and Wang et al. (2025b). These methods generate a specific token <search> that indicates an external search engine query is required. Once the retriever provides results, these are then fed back into the model as input, allowing it to resume content generation. Although these methods improved LLM utilization of retrievers, they do not truly provide a step-by-step approach for solving complex multi-hop problems. Their approaches are more concerned with how LLMs arrive at the correct answer, without ensuring the rigor, stability, or logical consistency of the underlying problem-solving process. Consequently, these methods are often unreliable in high-stakes or sensitive domains such as finance, medicine, and legal applications.
@@
In order to achieve logical reasoning with more accurate judgments and reduced biases (Li et al., 2025c), recent work has attempted various approaches, such as the atomic reasoning actions decomposition by Wu et al. (2024) , hierarchical and structured reasoning by Yang et al. (2025) , and verify reasoning trajectories with symbolic solver by Guan et al. (2025) . Moreover, Sumers et al. (2024) propose a cognitive architectures for language agents with three key dimensions: information storage, action space and decision-making procedure. Benefit from these previous works and inspired by Slow Thinking - which solves complex problem with access to additional computational resources, full attention, and sophisticated logical reasoning Booch et al. (2021) - we propose an innovative structured thinking and reasoning model as shown in Figure 1. It enables deep thinking in LLM by simulating human cognitive mechanisms without reinforcement learning DeepSeek-AI et al. (2025); OpenAI (2024). We use logical form to decompose complex problems broadly and solve them in depth. Logical form improves the logic, rigor, and stability of the LLM reasoning process. When humans solve complex problems, they only perform external information retrieval on knowledge that they do not know. Similarly to this process, we do not need to perform external knowledge retrieval for all decomposed sub-problems. Therefore, we design an LLM’s knowledge boundary determination module to reduce unnecessary external retrieval and mitigate noise from the retrieved content. If the LLM cannot confidently answer a certain sub-question using its own knowledge, we rely on external KBs. However, search results from external KB are often unreliable and noisy. To address this, we created a Focusing-and-Reasoning module to extract the core relevant information. To facilitate a clearer understanding, we will introduce these key innovations in detail in the following paragraphs.
@@
Logical Form based Reasoning. Prior RAG approaches Lewis et al. (2020); Trivedi et al. (2023); Li et al. (2025a), for both questions and sub-questions, are restricted to plain text representations, thereby preventing them from effectively utilizing high-quality structured knowledge. To address this issue, logical form proposes four operation functions tailored to frequently encountered questions (see the Appendix 8). Each operation comprises two components: a natural language segment (Step) and a logical function expression segment (Action). For generic retrievers like E5 and BGE-M3, we can directly employ the content within the logical form’s Step. Conversely, for structured knowledge retrievers, the Action portion of the logical form can be utilized. Crucially, logical form facilitates the propagation of dependencies among interdependent sub-problems. For example, consider "Step1: Who is the director of Hit Parade Of 1947? Action1: Retrieval(s=s1:film[‘Hit Parade Of 1947’], p=p1:director, o=o1:director)" followed by "Step2: When did #1 die? Action2: Retrieval(s=o1, p=p2:deathtime, o=o2:deathtime)". The variables and enable seamless transfer of text and expressions from the outcome of the first planning step. In this manner, the logical form enhances the logic, rigor, and stability of the LLM reasoning.
@@
Breadth Decomposition and Depth Solving. Direct retrieval often proves insufficient for complex multi-hop questions, failing to provide immediate answers. For example, "Which film has the director who died first, Hit Parade Of 1947 or Khiladi 420?". To address such issues, we have devised a model based on breadth decomposition and depth solving. At a macroscopic level, the decomposition of complex questions into logical forms is generally straightforward and typically does not necessitate recourse to external knowledge. For instance, the aforementioned example can be broken down into five constituent sub-questions: "Who is the director of Hit Parade Of 1947?", "When did #1 die?", "Who is the director of Khiladi 420?", "When did #3 die?", and "Which film was directed by the director who died first according to #2 and #4?". The resolution of each individual sub-problem, conversely, may entail ranging from to (maximum number of searches) retrieval operations, contingent upon its inherent complexity. This in-depth approach to sub-problem solving enables each sub-problem to assimilate adequate external knowledge, thereby reaching a solvable state.
@@
Knowledge Boundary Determination. Most prior approaches Asai et al. (2023); Gutierrez et al. (2024) initiate retrieval for every question or sub-question, even when the information resides within the LLM’s inherent knowledge. However, LLMs do not necessitate external lookups for such straightforward queries. To mitigate this redundancy, we design a knowledge boundary determination module that eliminates unnecessary external retrieval operations. Prior to knowledge boundary determination, the LLM first generates answers to the sub-problems. To ensure the precision of this knowledge boundary determination, we employ two strategies: (1) prompt-based confidence assessment and (2) likelihood-based confidence assessment. Only when both conditions are satisfied is considered that the solution to the sub-problem does not require external retrieval, and the answer generated by the LLM is directly adopted.
@@
Ultimately, our research yields a human-like structured iterative thinking and depth solving model offering substantial advancements, notably: (1) Facilitating more effective utilization of high-quality KBs, (2) Contributing to a more logically coherent, rigorous, and stable approach to complex problem solving, (3) Enabling autonomous decisions regarding the reliance on intrinsic model knowledge versus external information, thereby reducing unnecessary queries, and (4) Exhibiting a greater capacity to manage noise within retrieved data.
@@ 2 Related Work
Retrieval-Augmented Generation (RAG) integrates relevant, real-time information from external KBs through a retrieval mechanism, thereby enhancing its responses with contextual richness and factual grounding Borgeaud et al. (2022); Dong et al. (2023); Luo et al. (2024). Despite these advantages, a persistent challenge in RAG systems is that retrieved content can often be mixed with irrelevant or erroneous information due to the inherent limitations of current retrieval solutions. Moreover, predefined reasoning adopts structured and modular RAG pipelines that limit flexibility in evolving and open-ended tasks Li et al. (2025c, b).
To mitigate these challenges, researchers have introduced various components designed to enhance the robustness and effectiveness of RAG systems Gao et al. (2023); Li et al. (2024); Tan et al. (2024). A key focus in recent studies has been on enabling RAG systems to autonomously determine when to leverage retrieval mechanisms. For example, Jin et al. (2025b); Sun et al. (2025); Wang et al. (2025b) introduced a special token to explicitly trigger an external search, with the retrieval results subsequently integrated into the model to support content generation. However, Ren et al. (2025) identified limitations in LLM regarding their ability to accurately distinguish between internal and external knowledge, as well as their difficulty in reconciling conflicts between the two. To address this, they proposed Prior judgment and Poster judgments to evaluate confidence levels and assess response correctness. In parallel, another significant line of research has focused on developing adaptive strategies to optimize how retrieval is utilized, based on query complexity. Jeong et al. (2024) developed a classifier to intelligently decide between iterative, single, or no retrieval strategies depending on the complexity of the query. Building upon this idea, Islam et al. (2024) proposed a hybrid adaptive retrieval method: The model initially generates an answer without retrieval and then dynamically decides the need for retrieval based on its confidence, with the aim of balancing performance gains with inference speed. Furthermore, Tang et al. (2025) extended this work by applying a reinforcement learning-based framework to select the most appropriate retrieval strategy for a given query. While these diverse approaches have significantly advanced the integration and adaptive use of retrieval mechanisms within RAG systems, they currently lack a comprehensive, step-by-step framework for effectively addressing complex multi-hop reasoning problems. This critical area remains an open challenge for further research.
@@ 3 Approach
We propose an innovative complex problem solving model, as shown in Figure 2. At its core, this model uses logical form to enable both broad problem decomposition and in-depth problem solving. By integrating a knowledge boundary determination module, we minimize unnecessary external queries, balancing the LLM’s internal knowledge with external information. For sub-problems that require external information, a dedicated focusing and reasoning module is employed to reduce noise in the retrieved content and extract essential information.
@@ 3.1.2 Depth Solving
Constrained by sub-problem representations and retriever capabilities, many sub-problems defy resolution through single-turn retrievals. We have engineered a depth-oriented resolution strategy for Retrieval sub-problem, as detailed in the right part of Figure 2. During depth solving, we prompt the LLM to iteratively search across multi-level and multi-dimensional contexts. The iterative process continues until one of the following conditions is met: (1) the maximum number of turns is reached, or (2) the model generates a conclusive response enclosed between the designated answer tokens <answer> and </answer>. Before conducting depth solving of the Retrieval sub-problem, we first perform a knowledge boundary determination on the sub-problem to assess whether the LLM can directly answer it. If the LLM can confidently answer the current sub-problem, there is no need for depth solving, and the answer can be directly obtained; otherwise, an depth solving process will be initiated. After retrieving relevant content, we conduct focusing and reasoning analysis, allowing the model to determine contextually whether to initiate the next action. If a subsequent action is required, a new Logical Form is generated; otherwise, the sub-question answer is produced. This process iterates interactively until either the maximum number of turns is reached or the sub-question answer is obtained.
@@ 3.4 Logical Form Solver
Second, as shown in Figure 4, we build a hybrid graph retriever in KAG to retrieve this structured index. This hybrid retrieval method consists mainly of five steps: Step 1. DPR: Use the natural language Step part in the specific logical form to retrieve the top-k chunks by matching the name or content with vector similarity. Step 2. Seed Node Focusing: Identify initial relevant knowledge units, atomic queries, and entities of the sub-problem as seed nodes for Personalized PageRank (PPR). The key steps are shown in Figure 5. Step 3. PPR: The top-k entities, top-k knowledge units, and top-k atomic queries retrieved in Step 2 are integrated as seed nodes, and PPR Yang et al. (2024) runs on the entire graph. Identify the top-k chunks through probability propagation on the path. Step 4. Rerank: For the chunks obtained by DPR and PPR, top-k chunks are selected, respectively, and then rerank them through Reciprocal Rank Fusion (RRF) Cormack et al. (2009). Step 5. Collect Facts: The top-k chunks and the top-k SPOs matched by the logical form are sent to KAG-Thinker as supporting facts for answer generation in this step.
@@ 4.1 Experimental Settings
Evaluation Metric. While EM Yu et al. (2024) provides strict verification of answer precision; however,its binary nature is overly restrictive for open-ended scenarios where semantically equivalent answers may exist legal phrasing differences. The F1 metric therefore serves as a nuanced supplement, particularly valuable when evaluating free-form textual output. Comparison Methods. To systematically assess KAG-Thinker’s effectiveness, we establish a three-tiered evaluation model comprising the following state-of-the-art baselines: (1) Non-Retrieval Paradigms Naive Generation: Direct answer synthesis without external knowledge integration; Chain-of-Thought (CoT): Explicit cognitive pathway formalization through sequential reasoning traces Wei et al. (2022). (2) Retrieval-Augmented Architectures Naive RAG: Standard retrieval-generation pipeline without iterative optimization Lewis et al. (2020); IRCoT: Multi-cycle retrieval-reasoning coordination with dynamic feedback mechanisms Trivedi et al. (2023); Search-o1: Agent-mediated search workflow integration in reasoning processes Li et al. (2025a). (3) Reinforcement Learning Models R1-Gen: Pure RL-optimized generation without search engine interfacing DeepSeek-AI et al. (2025); Search-R1: Multimodal trajectory optimization combining search interactions with reasoning steps Jin et al. (2025b); ZeroSearch: Supervised LLM transformation into dual-function retrieval module (relevant/noisy document generation) Sun et al. (2025); StepSearch: PPO-based search LLM training with incremental exploration Wang et al. (2025b).